OpenStudy (anonymous):

i have one more problem this time not even sure how to start it solve each formula for the indicated letter 5x - 4y=8, for y anyone?

OpenStudy (anonymous):

hey, "solve for y" means get y by itself on one side of the equation, so you can start by subtracting 5x from both sides

OpenStudy (anonymous):

5x-5x-4y=8-5x -4y/4=8-5x / 4 am i right so far

OpenStudy (anonymous):

y=8- 5x / 4

OpenStudy (anonymous):

is that the correct answer they are after then ?